Question Three (a) Discuss the major causes of business failure (b) Westlands Ltd. has applied for a loan at Kay Cee Bec Bank. The company provided the bank [6 marks] with the following abbreviated portion of the financial information: Total assets Earnings before Interest and Tax Networking capital Book value of equity Accumulated retained earnings Total liabilities Sh. 163,000,000 7,900,000 4,200,000 18,000,000 16,000,000 57,000,000 Required: Based on Edward Altmans Z-score, should the bank approve Westlands Ltd. loan application. Give reason(s). . [19 marks]
Answer #1

1.

Major cause of business failure:

Lack of experience

Insufficient capital

locational disadvantage

inefficient inventory management.

Poor credit policies

2.

Z Score = 3.25 + 6.56(X1) + 3.26(X2) + 6.72(X3) + 1.05(X4)

where

X1 = Working Capital / Total Assets X2 = Retained Earnings / Total Assets

X3 = Earnings Before Interest & Tax / Total Assets

X4 = Total Book Equity /Total Liabilities

Z=3.25+6.56*4200000/163000000+3.26*16000000/163000000+6.72*7900000/163000000+1.05*18000000/57000000=4.40

This is not a safe score (less than 4.95) and hence should be not given loan